Spiderman, whose mass is 72.0 kg, is dangling on the free end of a 11.4-m-long rope, the other end of which is fixed to a tree limb above. By repeatedly bending at the waist, he is able to get the rope in motion, eventually getting it to swing enough that he can reach a ledge when the rope makes a ? = 60.8
Answer #1

work done is maximum change in potential energy =

mg L (1-cos A)

=72*9.8*11.4(1-0.4878)

=4123.78 joules
Answer #2

Work done here equals the change in potential energy, so does not matter what reference you take to measure the change in distance....

Work done = m*g*H ... where H is the change in height .. you dont need an absolute height in this case...

Work done = m*g* (R - R cos(60.8)) ... that's it...

you can also verifi this...assume that the end of rope where batman hangs initially is at a distance D1 from the ground...then the the distance after swinging is D1 + (r- r cos(60.8))..

so change in height is D1 + (R- Rcos(60)) - D1 = (R- Rcos(60.8))
Answer #3

work done = mgl*(1-cos60.8) =72*9.8*11.4*(1-cos60.8) =4119.5 J
Answer #4

work done by the gravitational force= -mgL(1-cos(60.8))

=-72*9.81*11.4*(1-0.4878)

= -4124.46 J